# FlexEdit: Marrying Free-Shape Masks to VLLM for Flexible Image Editing

### v1

@@ Abstract
Combining Vision Large Language Models (VLLMs) with diffusion models offers a powerful method for executing image editing tasks based on human language instructions. However, language instructions alone often fall short in accurately conveying user requirements, particularly when users want to add, replace elements in specific areas of an image. Luckily, masks can effectively indicate the exact locations or elements to be edited, while they require users to precisely draw the shapes at the desired locations, which is highly user-unfriendly. To address this, we propose FlexEdit, an end-to-end image editing method that leverages both free-shape masks and language instructions for Flexible Editing. Our approach employs a VLLM in comprehending the image content, mask, and user instructions. Additionally, we introduce the Mask Enhance Adapter (MEA) that fuses the embeddings of the VLLM with the image data, ensuring a seamless integration of mask information and model output embeddings. Furthermore, we construct FSMI-Edit, a benchmark specifically tailored for free-shape mask, including 8 types of free-shape mask. Extensive experiments show that our method achieves state-of-the-art (SOTA) performance in LLM-based image editing, and our simple prompting technique stands out in its effectiveness.The code and data can be found at https://github.com/A-new-b/flex˙edit.
@@ 1 Introduction
Image editing has significantly benefited from the exceptional generative capabilities of diffusion models (Li, Singh, and Grover 2023; Chen et al. 2023b; Xuan et al. 2024; Dhariwal and Nichol 2021; Liu, Ding, and Jiang 2023; Rombach et al. 2022a; Geng et al. 2024; Brooks, Holynski, and Efros 2023; Ye et al. 2023; Pan et al. 2024; Zhang et al. 2024). Trained on large scale datasets of text-image pairs, these models can generate high-quality images that align with text prompts. Instruction-based image editing extends these capabilities by using human language instructions to modify specific elements within an original image. However, researchers have found that language instructions can often be ambiguous and complex (Lai et al. 2024; Pi et al. 2023; Huang et al. 2024), thus the text encoders (Radford et al. 2021) fail to accurately represent this information, resulting in subpar editing outcomes. Large Language Models (LLMs) (Lyu et al. 2019; Ouyang et al. 2022), with their strong instruction-following capabilities, can understand complex human instructions and reason out clear and simple editing directives. Models like SmartEdit (Huang et al. 2024) have pioneered the use of LLMs as text encoders to comprehend complex instructions, allowing users to input complex commands for image editing. However, when it comes to scenarios that are difficult to describe with language, such as specifying precise locations within an image, user have no idea how to prompt the models. Even if the model can understand, such prompts are clearly not user-friendly (e.g., “$3^{th}$ row and $3^{th}$ column of the grid”).
@@
Users spend excessive time crafting these prompts, and the model may still fail to grasp their meaning, leading to unsatisfactory results (see Upper Figure 1). To address these challenges, visual prompts, such as masks, can effectively and clearly indicate the exact locations or elements to be edited, overcoming the limitations of language commands. However, most existing methods that utilize mask references require users to precisely draw the shapes at the desired locations, which is highly user-unfriendly and can be a barrier to effective use (Chen et al. 2023b). If the mask shapes are irregular or do not completely overlap with the objects, they often produce very poor results (see Lower Figure 1).
In real-world scenarios, the shapes users draw for masks are often random and may cover only a small part of the intended area. These free-shape masks challenge the model’s ability to understand user requirements. To address this issue, we created the Free Shape Mask Instruction Edit (FSIM-Edit) benchmark, which includes 80 and 125 single images and multiple images respectively, 8 types of free-shape mask. Evaluating existing models with this benchmark revealed that they perform poorly in free-shape mask setting, even after the fine-tuning.
@@
To overcome these issues, we modified the mask functionality in MagicBrush (Zhang et al. 2024), integrating complex instructions with image editingin the FSIM-Edit dataset and proposed FlexEdit, an innovative end-to-end image editing method that leverages both free-shape masks and language instructions for Flexible Editing (FlexEdit). FlexEdit employs a Vision Large Language Model (VLLM) to comprehend the image content, visual prompts, and user instructions. We introduced an adapter structure that fuses the mask represented embeddings of the VLLM with the image data, ensuring a seamless integration of visual information and model output embeddings. This integration enhances the model’s ability to understand and execute complex editing tasks accurately. Experiments demonstrate that FlexEdit significantly outperforms other methods under free-shape mask conditions. Notably, our simple prompting technique proves to be highly effective, setting a new standard for ease of use and precision in the field. FlexEdit represents a significant step forward in the realm of intelligent image editing, combining the strengths of VLLMs and diffusion models to provide a user-friendly and highly capable editing solution.
Our contributions can be summarized as follows:
@@ 2.1 Image Editing with Diffusion Models
Image editing has achieved exceptional performance thanks to the generative abilities of diffusion models (Brooks, Holynski, and Efros 2023; Zhang et al. 2024; Geng et al. 2024; Chen et al. 2023b; Pan et al. 2024; Ye et al. 2023). Brooks, Holynski, and Efros (2023) introduced InstructPix2Pix, an instruction-based image editing approach that modifies images based on human language instructions. To better align with human instructions, Zhang et al. (2024) proposed MagicBrush, extending text-only instructions to scenarios where a visual mask is provided. On the other hand, subject-driven image editing tasks aim to generate images conditioned on a customized subject and a text prompt that describes the context. For example, Anydoor (Chen et al. 2023b) can teleport target objects to new scenes at user-specified locations (mask) with desired shapes. Pan et al. (2024) proposed Largen for image inpainting that enables seamless inpainting of masked scene images, incorporating both the textual prompts and specified subjects. Singh et al. (2024) and Xuan et al. (2024) start discussing to use coarse visual prompt such as bounding boxes, scribble in guiding the editing, but they use the classic text encoder in diffusion models, which makes it hard to understand those complex instructions.
These models can combine user-provided images with visual prompts, such as masks, to perform image editing. The masks they use need to precisely overlap with the area to be edited, which is not user-friendly. In comparison, we propose using a free-shape mask as a visual prompt and introduce vision large language models (VLLMs) to better understand complex scenarios and instructions for visual prompt-based image editing in an end to end manner.
@@ 2.2 VLLMs in Image Editing
Vision Large Language Models (VLLMs) are known for their extensive world knowledge and ability to understand complex instructions (Bai et al. 2023; Lin et al. 2023; Liu et al. 2024, 2023; Li et al. 2023b; Chen et al. 2023a). Fine-tuning these VLLMs for image generation tasks has shown great success in comprehending human language instructions and generating images (Huang et al. 2024; Ge et al. 2024; Li, Singh, and Grover 2023). For example, Huang et al. (2024) fine-tune the LLaVA (Liu et al. 2024) model with dataset of MagicBrush (Zhang et al. 2024) to fit it into the image editing domain. While InstructAny2Pix (Li, Singh, and Grover 2023) build a multi-modal editing system that enables users to edit an image using instructions involving audio, images, and text.
Despite their success, current models predominantly rely on language-based instructions, overlooking the power of visual prompts such as masks. This reliance can limit the intuitiveness and accessibility of the editing process. Our approach seeks to enhance the editing capabilities by incorporating both user instructions and visual prompts, thereby achieving more precise and user-friendly image edits.
@@ 3 Methodology
This paper presents FlexEdit designed to integrate human instructions and visual prompts for complex image editing tasks. Existing methods, whether based on language instructions or subject mask editing, struggle with understanding user-provided free-shape masks (Chen et al. 2023b; Huang et al. 2024). This challenge lead to a situation that user struggle to draw a mask to meet the models’ need, which is not user-friendly and not acceptable.
@@ 3.1 FlexEdit Framework
We expanded the vocabulary of LLaVA (Liu et al. 2024) with $N$ new tokens “<$img0$>, …, <$imgN$>” to represent image editing information. As shown in Figure 3, $x_{1}$, $x_{2}$, $m$, and $\tau$ are input into the Vision Large Language Model (VLLM) $V(\cdot;\omega)$ to obtain response tokens $r_{i}$. Following this, the response set $R=\{r_{1},r_{2},\dots,r_{n}\}$ is processed to extract the corresponding hidden states $h_{i}\in\mathcal{H}$, capturing the deeper semantic information from the VLLM’s understanding of the inputs. However, these hidden states primarily reside in the LLM’s text vector space, which poses compatibility challenges when interfacing with the diffusion model-particularly one trained on CLIP (Radford et al. 2021) text embeddings. To address this, we introduce a Q-Former (Li et al. 2023c) module that refines the hidden states into embeddings $e$ compatible with the diffusion model. The transformation process is summarized as follows:
@@
For efficient training, we utilize LoRA (Hu et al. 2021) fine-tuning techniques, freezing most VLLM parameters. Let the ground truth text label of the VLLM response be $R^{\prime}$. The VLLM optimization process follows the recurrent loss calculation:
@@
During the diffusion process, the encoded image latent $z=\mathcal{E}(x_{1})$ is concatenated with the noisy latent $z_{t}$ and fed into the UNet (Ronneberger, Fischer, and Brox 2015) in a residual manner. The UNet $\epsilon_{\alpha}$ is trained to predict the noise added to the noisy latent $z_{t}$, with noise levels increasing under timesteps $t$, given the condition $c$ from the MEA module. The diffusion optimization process is formulated as:
@@ 3.2 Mask Enhanced Adapter
The scene $x_{1}$ and subject $x_{2}$ images are first mapped to the text embedding space via the CLIP image encoder, for CLIP’s training objective is to maximize the similarity of the image and text label embedding. The output $e$ of QFormer contains the editing intent generated by VLLM after combining the mask and instruction. However, it lacks fine-grained details from the original image. To reintroduce this information, we employed a cross-attention mechanism to fuse the output $e$ from QFormer (Li et al. 2023c) with the features from the scene image $x_{1}$, resulting the scene features $f_{1}$. The key and value of the cross-attention come from the scene image $x_{1}$ with the the output $e$ from QFormer as the query. Formulated as follow:
@@
Figure 5: Qualitative results for free-shape mask guided image editing in multiple image setting. We compare the subject image with no background and complex background.
Figure 6: Qualitative results comparison on single image edit.
3.3 Training to Understand Free-Shape Mask
@@
For easy annotation and efficient training, we modified the classical segmentation dataset COCO (Yu et al. 2016) and GRefCOCO (Liu, Ding, and Jiang 2023) for free-shape mask understanding training process. Specifically, we first adapt the original mask label $m_{o}$ into free-shape mask $m=W(m_{o})$ by a simple random walking algorithm $W$, and model are trained to predict the original mask while output the label of the predict class $p$. For example, the output can be “The masks means ‘chair’, <$img0\dots N$>”, <$img0\dots N$> represent the generate image with original mask $m_{o}$ (see Figure 4). More technical details will be discussed in the supplementary material.
3.4 FSMI-Edit Benchmark
We construct the Free-Shape Mask Instruction Edit Benchmark (FSMI-Edit) for evaluating image editing method in free-shape mask understanding. We focus on daily indoor (e.g. Backpack, Kitchen) and outdoor (e.g. Animals, Ground) scenarios, for subject image we also consider two types of both simple (only subject provided) and complex (subject with noisy background) subject image. FSMI-Edit consists of 80 and 125 single images and multiple images editing picture respectively.
We carefully design the mask type so that it can mimic human input preference, thoroughly evaluating the performance of free-shape mask reference image editing. Example of the free-shape mask from FSMI-Edit is shown at Figure 7. The free-shape mask consists of 8 type total: circle, circle(open hole), rectangle, rectangle (open hole), triangle, triangle (open hole), irregular, irregular (open hole) mask, with different orientation. We hope more researchers will pay attention to the free-shape mask guided image editing from these perspectives, thereby fostering the user-friendly multi-modal instruction image editing methods.
@@ 4 Experiments
Table 1: Comparison in multiple image editing domain, we compare the results in full mask and free-shape mask setting, and report the scores in foreground and the whole image.
[…]
Table 2: Comparison in single image editing domain.
[…]
@@ 4.1 Experiments Setting
Training Datasets
We include training data from four categories: (1) segmentation datasets, which are modified from COCO and GRefCOCO; (2) image editing datasets, which are adapted from MagicBrush (Zhang et al. 2024), ReasonEdit (Huang et al. 2024), and STRAT (Zhang et al. 2022); (3) in-context visual question answering (VQA) datasets, specifically Mimic-it (Li et al. 2023a), due to its capability of handling multi-image inputs; (4) our self-constructed image editing dataset, which includes a total of 390 pairs of images with complex subject image that should be learned from understand the prompt (e.g. “Add the backpack in [pic2] to the white mask in [pic1]” see Figure 5).
To address the imprecision of user-drawn masks in real-world applications, we avoid using the actual masks directly as inputs. Instead, we simulate user input by generating free shape masks using random walks on the real masks. These generated masks are then reviewed and re-annotated by three volunteers to ensure they closely mimic real-life user inputs. For the segmentation datasets, the model’s objective is to regenerate masks based on the free-form input to match the ground truth shapes. Training on the VQA dataset ensures that the model retains its analytical and reasoning capabilities during fine-tuning. Additional technical details can be found in the supplementary materials.
Metrics

### v2

@@ Abstract
Combining Vision Large Language Models (VLLMs) with diffusion models offers a powerful method for executing image editing tasks based on human language instructions. However, language instructions alone often fall short in accurately conveying user requirements, particularly when users want to add, replace elements in specific areas of an image. Luckily, masks can effectively indicate the exact locations or elements to be edited, while they require users to precisely draw the shapes at the desired locations, which is highly user-unfriendly. To address this, we propose FlexEdit, an end-to-end image editing method that leverages both free-shape masks and language instructions for Flexible Editing. Our approach employs a VLLM in comprehending the image content, mask, and user instructions. Additionally, we introduce the Mask Enhance Adapter (MEA) that fuses the embeddings of the VLLM with the image data, ensuring a seamless integration of mask information and model output embeddings. Furthermore, we construct FSMI-Edit, a benchmark specifically tailored for free-shape mask, including 8 types of free-shape mask. Extensive experiments show that our method achieves state-of-the-art (SOTA) performance in LLM-based image editing, and our simple prompting technique stands out in its effectiveness. The code and data can be found at https://anonymous.4open.science/r/FlexEdit
1Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences
2Shenzhen Technology University
3Georgia Institute of Technology
4Carnegie Mellon University
@@ 1 Introduction
Image editing has significantly benefited from the exceptional generative capabilities of diffusion models [13, 4, 29, 5, 15, 25, 7, 2, 30, 22, 32]. Trained on large scale datasets of text-image pairs, these models can generate high-quality images that align with text prompts. Instruction-based image editing extends these capabilities by using human language instructions to modify specific elements within an original image. However, researchers have found that language instructions can often be ambiguous and complex [10, 23, 9], thus the text encoders [24] fail to accurately represent this information, resulting in subpar editing outcomes. Large Language Models (LLMs) [18, 21], with their strong instruction-following capabilities, can understand complex human instructions and reason out clear and simple editing directives. Models like SmartEdit [9] have pioneered the use of LLMs as text encoders to comprehend complex instructions, allowing users to input complex commands for image editing.
@@
However, when it comes to scenarios that are difficult to describe with language, such as specifying precise locations within an image, user have no idea how to prompt the models. Even if the model can understand, such prompts are clearly not user-friendly (e.g., “$3^{th}$ row and $3^{th}$ column of the grid”). Users spend excessive time crafting these prompts, and the model may still fail to grasp their meaning, leading to unsatisfactory results (see Upper Figure 1). To address these challenges, visual prompts, such as masks, can effectively and clearly indicate the exact locations or elements to be edited, overcoming the limitations of language commands. The challenge of reducing the "tedious" and "time-consuming" effort of manual segmentation is well-documented in HCI literature, with foundational works like Intelligent Scissors [19] and GrabCut [27] created to simplify the process. Our work continues this user-centric philosophy. However, most contemporary methods that utilize mask references still require users to precisely draw shapes at the desired locations, which is highly user-unfriendly and can be a barrier to effective use [4]. If the mask shapes are irregular or do not completely overlap with the objects, they often produce very poor results (see Lower Figure 1).
@@
In real-world scenarios, the shapes users draw for masks are often random and may cover only a small part of the intended area. These free-shape masks challenge the model’s ability to understand user requirements. To address this issue, we created the Free Shape Mask Instruction Edit (FSIM-Edit) benchmark, which includes 80 and 125 single images and multiple images respectively, 8 types of free-shape mask. Evaluating existing models with this benchmark revealed that they perform poorly in free-shape mask setting, even after the fine-tuning. To overcome these issues, we modified the mask functionality in MagicBrush [32], integrating complex instructions with image editingin the FSIM-Edit dataset and proposed FlexEdit, an innovative end-to-end image editing method that leverages both free-shape masks and language instructions for Flexible Editing (FlexEdit). FlexEdit employs a Vision Large Language Model (VLLM) to comprehend the image content, visual prompts, and user instructions. We introduced an adapter structure that fuses the mask represented embeddings of the VLLM with the image data, ensuring a seamless integration of visual information and model output embeddings. This integration enhances the model’s ability to understand and execute complex editing tasks accurately. Experiments demonstrate that FlexEdit significantly outperforms other methods under free-shape mask conditions. Notably, our simple prompting technique proves to be highly effective, setting a new standard for ease of use and precision in the field. FlexEdit represents a significant step forward in the realm of intelligent image editing, combining the strengths of VLLMs and diffusion models to provide a user-friendly and highly capable editing solution. Our contributions can be summarized as follows:
@@ 2.1 Image Editing with Diffusion Models
While our work is situated within the paradigm of VLLM-guided diffusion models, it is worth noting the contributions of Generative Adversarial Networks (GANs) to image editing. GAN-based methods often offer faster inference speeds. However, diffusion models generally provide superior training stability, mode coverage, and higher generation fidelity, which are critical for the complex, instruction-guided editing tasks that we target. Therefore, we focus on diffusion models and their direct state-of-the-art competitors for the most relevant and fair comparison.
Image editing has achieved exceptional performance thanks to the generative abilities of diffusion models [2, 32, 7, 4, 22, 30]. [2] introduced InstructPix2Pix, an instruction-based image editing approach that modifies images based on human language instructions. To better align with human instructions, [32] proposed MagicBrush, extending text-only instructions to scenarios where a visual mask is provided. On the other hand, subject-driven image editing tasks aim to generate images conditioned on a customized subject and a text prompt that describes the context. For example, Anydoor [4] can teleport target objects to new scenes at user-specified locations (mask) with desired shapes. [22] proposed Largen for image inpainting that enables seamless inpainting of masked scene images, incorporating both the textual prompts and specified subjects. [28] and [29] start discussing to use coarse visual prompt such as bounding boxes, scribble in guiding the editing, but they use the classic text encoder in diffusion models, which makes it hard to understand those complex instructions. These models can combine user-provided images with visual prompts, such as masks, to perform image editing. The masks they use need to precisely overlap with the area to be edited, which is not user-friendly. In comparison, we propose using a free-shape mask as a visual prompt and introduce vision large language models (VLLMs) to better understand complex scenarios and instructions for visual prompt-based image editing in an end to end manner.
@@ 2.2 VLLMs in Image Editing
Vision Large Language Models (VLLMs) are known for their extensive world knowledge and ability to understand complex instructions [1, 14, 17, 16, 11, 3]. Fine-tuning these VLLMs for image generation tasks has shown great success in comprehending human language instructions and generating images [9, 6, 13]. For example, [9] fine-tune the LLaVA [17] model with dataset of MagicBrush [32] to fit it into the image editing domain. While InstructAny2Pix [13] build a multi-modal editing system that enables users to edit an image using instructions involving audio, images, and text. Despite their success, current models predominantly rely on language-based instructions, overlooking the power of visual prompts such as masks. This reliance can limit the intuitiveness and accessibility of the editing process. Our approach seeks to enhance the editing capabilities by incorporating both user instructions and visual prompts, thereby achieving more precise and user-friendly image edits.
@@ 3 Methodology
This paper presents FlexEdit designed to integrate human instructions and visual prompts for complex image editing tasks. Existing methods, whether based on language instructions or subject mask editing, struggle with understanding user-provided free-shape masks [4, 9]. This challenge lead to a situation that user struggle to draw a mask to meet the models’ need, which is not user-friendly and not acceptable.
@@ 3.1 FlexEdit Framework
We expanded the vocabulary of LLaVA [17] with $N$ new tokens “<$img0$>, …, <$imgN$>” to represent image editing information. As shown in Figure 3, $x_{1}$, $x_{2}$, $m$, and $\tau$ are input into the Vision Large Language Model (VLLM) $V(\cdot;\omega)$ to obtain response tokens $r_{i}$. Following this, the response set $R=\{r_{1},r_{2},\dots,r_{n}\}$ is processed to extract the corresponding hidden states $h_{i}\in\mathcal{H}$, capturing the deeper semantic information from the VLLM’s understanding of the inputs. However, these hidden states primarily reside in the LLM’s text vector space, which poses compatibility challenges when interfacing with the diffusion model-particularly one trained on CLIP [24] text embeddings. To address this, we introduce a Q-Former [12] module that refines the hidden states into embeddings $e$ compatible with the diffusion model. The transformation process is summarized as follows:
@@
For efficient training, we utilize LoRA [8] fine-tuning techniques, freezing most VLLM parameters. Let the ground truth text label of the VLLM response be $R^{\prime}$. The VLLM optimization process follows the recurrent loss calculation:
@@
During the diffusion process, the encoded image latent $z=\mathcal{E}(x_{1})$ is concatenated with the noisy latent $z_{t}$ and fed into the UNet [26] in a residual manner. The UNet $\epsilon_{\alpha}$ is trained to predict the noise added to the noisy latent $z_{t}$, with noise levels increasing under timesteps $t$, given the condition $c$ from the MEA module. The diffusion optimization process is formulated as:
@@ 3.2 Mask Enhanced Adapter
The scene $x_{1}$ and subject $x_{2}$ images are first mapped to the text embedding space via the CLIP image encoder, for CLIP’s training objective is to maximize the similarity of the image and text label embedding. The output $e$ of QFormer contains the editing intent generated by VLLM after combining the mask and instruction. However, it lacks fine-grained details from the original image. To reintroduce this information, we employed a cross-attention mechanism to fuse the output $e$ from QFormer [12] with the features from the scene image $x_{1}$, resulting the scene features $f_{1}$. The key and value of the cross-attention come from the scene image $x_{1}$ with the the output $e$ from QFormer as the query. Formulated as follow:
@@
3.3 Training to Understand Free-Shape Mask
For easy annotation and efficient training, we modified the classical segmentation dataset COCO [31] and GRefCOCO [15] for free-shape mask understanding training process. Specifically, we first adapt the original mask label $m_{o}$ into free-shape mask $m=W(m_{o})$ by a simple random walking algorithm $W$, and model are trained to predict the original mask while output the label of the predict class $p$. For example, the output can be “The masks means ‘chair’, <$img0\dots N$>”, <$img0\dots N$> represent the generate image with original mask $m_{o}$ (see Figure 4). More technical details will be discussed in the supplementary material.
@@ 3.3 Training to Understand Free-Shape Mask
3.4 FSMI-Edit Benchmark
To address the lack of standardized evaluation for this task, we construct the Free-Shape Mask Instruction Edit (FSMI-Edit) benchmark. To our knowledge, it is the first benchmark specifically designed to evaluate image editing under the realistic condition of imprecise, free-shape masks. We focus on daily indoor (e.g. Backpack, Kitchen) and outdoor (e.g. Animals, Ground) scenarios, for subject image we also consider two types of both simple (only subject provided) and complex (subject with noisy background) subject image. FSMI-Edit consists of 80 and 125 single images and multiple images editing picture respectively. We carefully design the mask type so that it can mimic human input preference, thoroughly evaluating the performance of free-shape mask reference image editing. Example of the free-shape mask from FSMI-Edit is shown at Figure 5. The free-shape mask consists of 8 type total: circle, circle(open hole), rectangle, rectangle (open hole), triangle, triangle (open hole), irregular, irregular (open hole) mask, with different orientation. We hope more researchers will pay attention to the free-shape mask guided image editing from these perspectives, thereby fostering the user-friendly multi-modal instruction image editing methods.
@@ 4 Experiments
Table 1: Comparison in multiple image editing domain, we compare the results in full mask and free-shape mask setting.
Method Foreground Whole Image PNSR $\uparrow$ CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINOv2 $\uparrow$ PNSR $\uparrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ DINOv2 $\uparrow$ Full Mask AnyDoor 13.1316 22.0478 0.8959 0.2485 19.2274 0.1055 0.2340 0.9292 Largen 12.5130 20.8105 0.8332 0.2468 18.6747 0.1166 0.2331 0.8886 InstructAny2Pix 8.6129 20.9438 0.7557 0.2424 10.3434 0.5231 0.1723 0.7100 FlexEdit 13.1645 24.1754 0.8279 0.2982 18.0522 0.1433 0.2402 0.8596 Free-Shape Mask AnyDoor 11.8483 19.5245 0.7711 0.2451 18.4441 0.1348 0.2280 0.8292 Largen 12.1133 19.1737 0.7627 0.2386 18.6563 0.1373 0.2266 0.8271 InstructAny2Pix 9.6056 20.5750 0.7582 0.2411 11.4099 0.4720 0.1775 0.7151 FlexEdit 12.5648 22.4813 0.7795 0.2908 17.9068 0.1585 0.2353 0.8262
@@ 4.1 Experiments Setting
4.1.1 Training Datasets
Figure 6: Qualitative results for free-shape mask guided image editing in multiple image setting.
We include training data from four categories, with a large-scale, mixed composition: (1) 65To address the imprecision of user-drawn masks in real-world applications, we avoid using the actual masks directly as inputs. Instead, we simulate user input by generating free shape masks using random walks on the real masks. These generated masks are then reviewed and re-annotated by three volunteers to ensure they closely mimic real-life user inputs. For the segmentation datasets, the model’s objective is to regenerate masks based on the free-form input to match the ground truth shapes. Training on the VQA dataset ensures that the model retains its analytical and reasoning capabilities during fine-tuning. Additional technical details can be found in the supplementary materials.
4.1.2 Metrics